OpenStudy (anonymous):

Solve the equation. 4x^3-x=0

OpenStudy (anonymous):

don't understand..

OpenStudy (anonymous):

never mind i got it i think the answer is 0, 1/2 and -1/2

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

factorize it x(4x^2-1)=0 u will get x=0 or x^2=1/4=>x=+/-1/2